# A prime number “game of life”: can $\lfloor y\cdot p\#\rfloor$ be prime for all $p\geq 2$?

### v1

@@ 2. Definitions
$p\#$ the primorial function: $p\#=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot\ldots\cdot p$
@@
$\Phi(n,p)$ number of integers $\leq n$ that are relatively prime to $p\#$
@@ 3. Setting the stage
Lemma: Any real number $y$ such that $q=\lfloor y\cdot p\#\rfloor$ is a prime number for every $p\geq 2$ is contained in the interval [1,1.5[[1,1.5[[ 1 , 1.5 [.
@@
For $y\geq 1.5$, $\lfloor y\cdot 2\#\rfloor$ must be an odd prime $12m-$ 1 for some integer $m>1$. Then $\lfloor y\cdot 3\#\rfloor$ is a number in the interval $[6m-3,6m-1]$, only the latter of which can be prime because $36m-$ 3 is divisible by 3 and $26m-$ 2 is divisible by 2. Likewise, any integer in the successive intervals $[p\#\cdot m-p,p\#\cdot m-1]$ is divisible by a small prime $\leq p$ except for $p\#\cdot m-1$.
Hence there would have to be an integer $m$ for which $p\#\cdot m-1$ is always prime, i.e. the sequence $f(0)=m-1$, $f(s+1)=(f(s)+1)\cdot p_{s}-1$ must only produce primes for all $s\geq 1$, which is practically impossible (though we have not seen this being explicitly disproven yet!).
@@
[…]
@@
Table 2. This seems to be the new “game of life”, a family tree of prime numbers. Prime triplets $q+\{0,2,6\}$ or $q+\{0,4,6\}$ like in stage 8 are rather rare in the process, one finds the next example at stage 152 ($p$ = 881), and there are no prime quadruplets $q+\{0,2,6,8\}$ in any stage $\leq$ 318.
@@ 4. Analyzing the game
The size of the primes at each stage $s$ with the corresponding $p$, $p_{s}$ being the $s^{th}$ prime, is always about $1.25\cdot p_{s}\#$. The probability of a random number of this size being prime is $1/(\log p_{s}\#+0.23)$, or $1/(\theta(p_{s})+0.23)$ where $\theta(p)$ is the first Chebyshev function. For one prime $q$ of the stage $s-1$ there has to be on average at least one prime in the interval $[q\cdot p_{s},(q+1)\cdot p_{s}]$ to have a chance that the sequence keeps on producing ever more primes. More precisely, the probability of getting at least one prime out of the respective interval must be bigger than getting no prime at all.
@@
Table 4. For a random interval start and interval length 10, the average of numbers that are coprime to $2310=11\#$ is above two, while for an interval of the same length after a prime number (or any number coprime to $11\#$), significantly fewer than two such numbers are to be expected. Hardy and Littlewood [3] laid a solid foundation regarding this issue.
@@
prime +10 0% 50% 0% 17% 10% 38% 1.78
Ultimately, the number of numbers that are relatively prime to $k\#$ in the intervals in question $[q\cdot p+1,(q+1)\cdot p-1]$ — denoted by $\Phi(p-1,k)$ — is
$\begin{split}\prod\limits_{\begin{subarray}{c}u=3\\ u\,prime\end{subarray}}^{k}(1-\frac{1}{u-1})\cdot\bigg{(}\lfloor\frac{p-1}{2}\rfloor+\sum\limits_{\begin{subarray}{c}v=3\\ v\,prime\end{subarray}}^{\min(\frac{p-1}{2},k)}\frac{\lfloor\frac{p-1}{2v}\rfloor}{v-2}+\sum\limits_{\begin{subarray}{c}v_{1}=3\\ v_{1}\,prime\end{subarray}}^{\min(\sqrt{\frac{p-1}{2}},k)}\sum\limits_{\begin{subarray}{c}v_{2}=v_{1}+2\\ v_{2}\,prime\end{subarray}}^{\min(\frac{p-1}{2v_{1}},k)}\frac{\lfloor\frac{p-1}{2v_{1}v_{2}}\rfloor}{(v_{1}-2)(v_{2}-2)}\\ +\sum\limits_{\begin{subarray}{c}v_{1}=3\\ v_{1}\,prime\end{subarray}}^{\min(\sqrt[3]{\frac{p-1}{2}},k)}\sum\limits_{\begin{subarray}{c}v_{2}=v_{1}+2\\ v_{2}\,prime\end{subarray}}^{\min(\sqrt{\frac{p-1}{2v_{1}}},k)}\sum\limits_{\begin{subarray}{c}v_{3}=v_{2}+2\\ v_{3}\,prime\end{subarray}}^{\min(\frac{p-1}{2v_{1}v_{2}},k)}\frac{\lfloor\frac{p-1}{2v_{1}v_{2}}\rfloor}{(v_{1}-2)(v_{2}-2)(v_{3}-2)}+...\bigg{)}.\end{split}$ (5)
@@
As $k\to\infty$, the adjusted interval length (for our case with the semiprimes) is — apart from $p=3,7,\text{ and }13$ — always a bit smaller than $p-1$. Dividing it again by $\log(q\cdot p)$ and we arrive at the expected average number of primes in one interval. The ratio of the adjusted interval length vs. $p-1$, which is equivalent to $\frac{\Phi(p-1,k)}{W(k)\cdot(p-2)}$ for appropriately large $k$, will be denoted below by $\psi(p)$.
@@ 5. Proposing practical predictions
Again, we should bear in mind that these primes are not as flexible as (8) takes them to be: the probability of obtaining $n$ primes in the following step is zero when $n$ exceeds the maximal possible number of primes in the given interval (e.g. $n>97$ for $p-1=546$) as fixed in the specific rules for the k-tuple conjecture. To take account of that — at least to some degree —, $W(k)$ from (6) is deployed again such that $k\#<p$. For $p=547$, we set $k=7$ and $W(k)$ = 48/210 = 1/4.375. As $p$ gets larger than 11# = 2310, $W(k)$ can be changed to 480/2310 = 1/4.8125 and so on.
@@
with error terms depending on either $p$ or $\omega(s)$. As long as $m\approx-1$ as is expected on average in the long run, if $\omega(s)$ is chosen a bit larger than $1-1/\sqrt{2p_{s}}$, the values for the subsequent $\omega(s+x)$ by the recursion formula above quickly head toward 1, so the first term in the parenthesis above, $m/\sqrt{p}$, becomes the most significant. If then at some point $Li(p)\approx\pi(p)$, thus $m\approx 0$, while $\omega(s)$ is very close to one, then the second term 1/$p$ becomes the most significant, but doesn’t nearly have the (then opposite) impact on $\omega(s)$ as in the region where $m\approx-1$. Now what does that say? If $\omega(s)$ stays close to 1, the initially chosen value is larger than the actual value, which is good because the actual value — the probability that the sequence terminates — is desired to be as small as possible. An $m$-value of +1 would adequately counter the effect of the expected average value $-1$, then again that should happen just as often as $m\leq-3$. It should be noted that there is yet no known effective region where $|m+1|\geq 2$, making it a candidate for a future endeavor. 222 Myerscough [9] tabulates the logarithmic density for the case $m>1$ as $1.603\cdot 10^{-95}$.
@@
Most of the time, the actual value lies between $n-\sqrt{n}$ and $n+\sqrt{n}$, and, by standard deviation measures, more than 99.7% of the time it should be well within $n\pm 3\sqrt{n}$. But even considering a negative deviation of $4\sqrt{n}$: $-4\cdot\sqrt{20298}=-569.88...$ is still outnumbered by the predicted growth $20298-19690=608$, corroborating that the initial assumption holds with a very high probability if the distribution is normal (a sufficient but not particularly necessary condition).
@@
which is surprisingly accurate in spite of its simplicity in the narrow sense that it is off by less than 20, with one little exception, for $s\leq 61$, off by at most 20% for $98\leq s\leq 114$, and off by at most 15.4% for $115\leq s\leq 318$ — compared to more elaborate predictions, the latter may hold even for a couple more $s\leq 400$. Furthermore, (11) may be off by a factor of no more than 2 for $66\leq s\leq 1000+$. In light of the time it would take to meticulously calculate $n$ for all $s\leq 1000$, the exact crossover on the large side of the inequality will always remain a mystery, thus calling into question any more daring predictions regarding (11).
@@
At stage 318 ($p=2111$) there are 592642 probable primes, and $\omega(318)\approx 95.408884\%$. $0.95408884^{592642}\approx 2.99\cdot 10^{-12097}$ — that is, for the time being, the conjectured probability that the game fails for some $s>318$.
Up to that point, 23036547 numbers have not yet been tested for certified primality, the smallest of those being in the ballpark of $2.38\cdot 10^{247}$. No counterexample is known for the combination of a Baillie-Pomerance-Selfridge-Wagstaff pseudoprimality test and a strong Lucas test which were carried out on the probable primes. 333 including some trial division, additionally decreasing the chances that a pseudoprime slipped through However if there is one among all the established PRPs, it could take effect on at most 61642 PRPs (the most prolific branch at $p=601$). $0.95408884^{592642-61642}\approx 4.59\cdot 10^{-10839}$, i.e. $1.54\cdot 10^{1258}$ times larger. It is yet unclear how much this will affect the above evaluation, especially by taking a look at Pomerance [13] who gives a heuristic argument that the number of counterexamples for a BPSW primality test up to $x$ is $\gg x^{1-\varepsilon}$ for some $\varepsilon>0$ and sufficiently large $x$. In either case, said ratio can be systematically reduced by checking the primality of the prolific branches in the sequence.
@@
Now, what about the initially proposed values for $y$? The lower and upper bounds are easily obtained by picking the smallest prime $q_{s,1}$ and the largest prime $q_{s,n}$ of the last computed stage and dividing by $p\#$: $y_{\min}=q_{s,1}/p\#$ and $y_{\max}=(q_{s,n}+1)/p\#$. The calculations so far demonstrate that any $y$ lies in a very small range between
@@ 8. Ancestors and survivors
Having calculated all primes up to a certain stage, we can trace back these primes for previous stages and see which of those have “survived” in the long run. All probable primes from stage 318 originated from one prime at stage 43. Recall that there were a total of 14 primes at stage 43, so the other 13 “petered out” along the way, including one branch that originated in stage 37 tenaciously keeping up until it succumbed at stage 95.
@@
Figure 3. Number of possible values for $y$ (for all we know). From $s=95$ onward, all primes $q$ emerge from one common ancestor at $s=43$, namely 1292942159746921794791923187781692727375711831825607985864285936838920812561. Naturally, the branches in the picture are not equally strong. For $s=318$ ($p=2111$), the following number of probable primes divide up on the nine given groups: 8080 + 7132 + 48289 + 90644 + 100189 + 74398 + 10066 + 56007 + 197837.
If one was to specify these values as $y_{0}$, $y_{1}$ and so on to attain fixed constants for the various possibilities, one might set $y_{0}$ as the minimum value $y_{min}$ as described above, and label each of the consecutive numbers $y_{1}$, $y_{2}$ as the other smallest possible value(s) after a precalculated split, when it is “safe enough to assume” that those branches are stable.
@@
(from the first split at stage 44, $\lfloor y_{0}\cdot 191\#\rfloor\cdot 193+30$ and $+88$, $y_{1}\approx y_{0}+58/193\#$)
@@
The value $y_{6}$ is one that may be doubted more than $y_{0}$$y_{5}$ since it derives from the weakest branch of Figure 3 with only 7132 primes. But since 7132 $>$ 2111 (thus a stable branch), it can be expected to remain valid.
@@
The bifurcation tree as depicted in Figure 3 gets quite impressive when expanded. For example, tracing back the primes from stage 318 to stage 100 crystallizes out 47 primes (compared to originally 594) with an eventful history.
@@ 9. Broader utilization
[…]
@@
$10<y<$ 1: this gives a sequence which is stronger than the original p#Y-sequence
($21<y<$ 2: original sequence)
$2625<y<$ 26: stronger than the original sequence up to stage 166, weaker after stage 168
$14121411<y<$ 1412: a weak sequence with only 883 primes at $s=168$, $p=997$
$34333432<y<$ 3433: a very weak sequence, only 168 primes at $s=168$, $p=997$
$1394913948<y<$ 13949: another weak sequence
$201421201420<y<$ 201421: another one which has somewhat good chances to survive
$60071046007103<y<$ 6007104: rather weak
@@
$32<y<$ 3: extinct at stage 4
$43<y<$ 4: extinct at stage 7
$65<y<$ 6: extinct at stage 11
$1716<y<$ 17: extinct at stage 14
$979978<y<$ 979: extinct at stage 17
$66416640<y<$ 6641: extinct at stage 23
$1145711456<y<$ 11457: extinct at stage 35
$160564160563<y<$ 160564: extinct at stage 38
$283258283257<y<$ 283258: extinct at stage 68
$11172311117230<y<$ 1117231: extinct at stage 78
$15945021594501<y<$ 1594502: extinct at stage 86
@@
Table 7. Data for $s\leq 318$, using BPSW-pseudoprimality for the primes $q$ for $s\geq 110$. $n*$ stands for the number of surviving branches, backtracked from the primes of stage 318. This number seems to be settled for $s\leq 130$ but will be subject to change for $s\geq 131$ by the time the calculation is taken further since many of them will get cancelled out, so this column only gives a momentary picture (as well as an upper bound). Nevertheless, the rate of decrease for larger s may be of interest in itself.
$s$ $p$ $n$ $s$ $p$ $n$ $n*$ $s$ $p$ $n$ $n*$ $s$ $p$ $n$ $n*$

### v2

@@ 2. Definitions
p#p\# the primorial function: $p\#=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot\ldots\cdot p$
@@
$\Phi(n,p)$ number of integers $\leq n$ that are relatively prime to p#p\#
@@ 3. Setting the stage
[…]
Lemma: Any real number $y$ such that $q=\lfloor y\cdot p\#\rfloor$ is a prime number for every $p\geq 2$ is contained in the interval $[1,1.5[$.
@@
For $y\geq 1.5$, $\lfloor y\cdot 2\#\rfloor$ must be an odd prime 2m12m-1 for some integer $m>1$. Then $\lfloor y\cdot 3\#\rfloor$ is a number in the interval [6m3,6m1][6m-3,6m-1], only the latter of which can be prime because 6m36m-3 is divisible by 3 and 6m26m-2 is divisible by 2. Likewise, any integer in the successive intervals $p\#\cdot m-p,p\#\cdot m-1]$ is divisible by a small prime $\leq p$ except for $p\#\cdot m-1$.
Hence there would have to be an integer $m$ for which $p\#\cdot m-1$ is always prime, i.e. the exponentially growing sequence f(0)=m1f(0)=m-1, $f(s+1)=(f(s)+1)\cdot p_{s}-1$ must only produce primes for all $s\geq 1$, which is practically impossible (though we have not seen this being explicitly disproven yet!).
@@
$\bullet\,$ Stage 1: $p=2\,\rightarrow\,2\cdot Y=[2,3[\,\rightarrow\,\lfloor 2\cdot Y\rfloor=2$ – no further action required yet
$\bullet\,$ Stage 2: $p=3\,\rightarrow\,6\cdot Y=[6,9[$ – the only prime in here is 7, so Y narrows down to $\frac{7}{6},\frac{8}{6}[$
$\bullet\,$ Stage 3: $p=5\,\rightarrow\,30\cdot Y=[35,40[$ – one prime (37) in this interval, $Y$ shrinks to $\frac{37}{30},\frac{38}{30}[$
$\bullet\,$ Stage 4: $p=7\,\rightarrow\,210\cdot Y=[259,266[$ – once again, one prime in here and any possible $y$ lies between $\frac{263}{210}$ and $\frac{264}{210}$
$\bullet\,$ Stage 5: $p=11\,\rightarrow\,2310\cdot Y=[2893,2904[$ – now it’s getting interesting, since there are two primes in that interval, 2897 and 2903. We will split $Y$ into two separate intervals, $Y_{1}=[\frac{2897}{2310},\frac{2898}{2310}[$ and $Y_{2}=[\frac{2903}{2310},\frac{2904}{2310}[$ and proceed
$\bullet\,$ Stage 6: $p=13\,\rightarrow\,$ each $Y_{1}$ and $Y_{2}$ brings forth one prime, 37663 and 37747 respectively. Cutting back both $Y$’s, then
$\bullet\,$ Stage 7: $p=17\,\rightarrow\,$ $Y_{1}$ leads to 640279, whereas $Y_{2}$ gives 641701 and 641713, and $Y_{2}$ will be split into $Y_{2}$ and $Y_{3}$ in accordance with the primes
$\bullet\,$ Stage 8: $p=19\,\rightarrow\,9699690\cdot Y_{1}$ contains a prime triplet, 12165311+$d$ for $d=\{0,2,6\}$, while $Y_{2}$ and $Y_{3}$ both fail to produce any further prime
$\bullet\,$ etc.
@@
Table 2. This seems to be the new “game of life”, a family tree of prime numbers. Prime triplets $q+\{0,2,6\}$ or $q+\{0,4,6\}$ like in stage 8 are rather rare in the process, one finds the next example at stage 152 ($p$ = 881), and there are no prime quadruplets $q+\{0,2,6,8\}$ in any stage $\leq$ 331.
@@ 4. Analyzing the game
The size of the primes at each stage $s$ with the corresponding $p$, $p_{s}$ being the $s^{th}$ prime, is always about $1.25\cdot p_{s}\#$. The probability of a random number of this size being prime is $1/(\log p_{s}\#+0.23)$, or $1/(\theta(p_{s})+0.23)$ where $\theta(p)$ is the first Chebyshev function. For one prime $q$ of the stage s1s-1 there has to be on average at least one prime in the interval $q\cdot p_{s},(q+1)\cdot p_{s}]$ to have a chance that the sequence keeps on producing ever more primes. More precisely, the probability of getting at least one prime out of the respective interval must be bigger than getting no prime at all.
@@
Table 4. For a random interval start and interval length 10, the average of numbers that are coprime to 2310=11#2310=11\# is above two, while for an interval of the same length after a prime number (or any number coprime to 11#11\#), significantly fewer than two such numbers are to be expected. Hardy and Littlewood [3] laid a solid foundation regarding this issue.
@@
[…]
Ultimately, the number of numbers that are relatively prime to k#k\# in the intervals in question $q\cdot p+1,(q+1)\cdot p-1]$ — denoted by $\Phi(p-1,k)$ — is
[…]
@@
As $k\to\infty$, the adjusted interval length (for our case with the semiprimes) is — apart from $p=3,7,\text{ and }13$ — always a bit smaller than p1p-1. Dividing it again by $\log(q\cdot p)$ and we arrive at the expected average number of primes in one interval. The ratio of the adjusted interval length vs. p1p-1, which is equivalent to $\frac{\Phi(p-1,k)}{W(k)\cdot(p-2)}$ for appropriately large $k$, will be denoted below by $\psi(p)$.
@@ 5. Proposing practical predictions
Again, we should bear in mind that these primes are not as flexible as (8) takes them to be: the probability of obtaining $n$ primes in the following step is zero when $n$ exceeds the maximal possible number of primes in the given interval (e.g. $n>97$ for p1=546p-1=546) as fixed in the specific rules for the k-tuple conjecture. To take account of that — at least to some degree —, $W(k)$ from (6) is deployed again such that k#<pk\#<p. For $p=547$, we set $k=7$ and $W(k)$ = 48/210 = 1/4.375. As $p$ gets larger than 11# = 2310, $W(k)$ can be changed to 480/2310 = 1/4.8125 and so on.
@@
with error terms depending on either $p$ or $\omega(s)$. As long as $m\approx-1$ as is expected on average in the long run, if $\omega(s)$ is chosen a bit larger than $1-1/\sqrt{2p_{s}}$, the values for the subsequent $\omega(s+x)$ by the recursion formula above quickly head toward 1, so the first term in the parenthesis above, $m/\sqrt{p}$, becomes the most significant. If then at some point $Li(p)\approx\pi(p)$, thus $m\approx 0$, while $\omega(s)$ is very close to one, then the second term 1/$p$ becomes the most significant, but doesn’t nearly have the (then opposite) impact on $\omega(s)$ as in the region where $m\approx-1$. Now what does that say? If $\omega(s)$ stays close to 1, the initially chosen value is larger than the actual value, which is good because the actual value — the probability that the sequence terminates — is desired to be as small as possible. An $m$-value of +1 would adequately counter the effect of the expected average value 1-1, then again that should happen just as often as $m\leq-3$. It should be noted that there is yet no known effective region where $|m+1|\geq 2$, making it a candidate for a future endeavor. 222 Myerscough [9] tabulates the logarithmic density for the case $m>1$ as $1.603\cdot 10^{-95}$.
@@
Most of the time, the actual value lies between $n-\sqrt{n}$ and $n+\sqrt{n}$, and, by standard deviation measures, more than 99.7% of the time it should be well within $n\pm 3\sqrt{n}$. But even considering a negative deviation of $4\sqrt{n}$: $4\cdot\sqrt{20298}=-569.88$... is still outnumbered by the predicted growth 2029819690=60820298-19690=608, corroborating that the initial assumption holds with a very high probability if the distribution is normal (a sufficient but not particularly necessary condition).
@@
which is surprisingly accurate in spite of its simplicity in the narrow sense that it is off by less than 20, with one little exception, for $s\leq 61$, off by at most 20% for $98\leq s\leq 114$, and off by at most 15.4% for $115\leq s\leq 331$ — compared to more elaborate predictions, the latter may hold even for a couple more $s\leq 500$. Furthermore, (11) may be off by a factor of no more than 2 for $66\leq s\leq 1000+$. In the long run, however, the expression in (11) is expected to far exceed $n$, thus calling into question any more daring predictions based on it.
@@
At stage 331 ($p=2221$) there are 810387 probable primes, and $\omega(331)\approx 95.45431\%$. $0.9545431^{810387}\approx 3.71\cdot 10^{-16374}$ — that is, for the time being, the conjectured probability that the game fails for some $s>331$.
Up to that point, 32103994 numbers have not yet been tested for certified primality, the smallest of those being in the ballpark of $2.38\cdot 10^{247}$. No counterexample is known for the combination of a Baillie-Pomerance-Selfridge-Wagstaff pseudoprimality test and a strong Lucas test which were carried out on the probable primes. 333 including some trial division, additionally decreasing the chances that a pseudoprime slipped through However if there is one among all the established PRPs, it could take effect on at most 85448 PRPs (the most prolific branch at $p=601$). $0.9545431^{810387-85448}\approx 1.00\cdot 10^{-14647}$, i.e. $2.70\cdot 10^{1726}$ times larger. It is yet unclear how much this will affect the above evaluation, especially by taking a look at Pomerance [13] who gives a heuristic argument that the number of counterexamples for a BPSW primality test up to $x$ is $\gg x^{1-\varepsilon}$ for some $\varepsilon>0$ and sufficiently large $x$. In either case, said ratio can be systematically reduced by checking the primality of the prolific branches in the sequence.
@@
Now, what about the initially proposed values for $y$? The lower and upper bounds are easily obtained by picking the smallest prime $q_{s,1}$ and the largest prime $q_{s,n}$ of the last computed stage and dividing by p#p\#: $y_{\min}=q_{s,1}/p\#$ and $y_{\max}=(q_{s,n}+1)/p\#$. The calculations so far demonstrate that any $y$ lies in a very small range between
@@ 8. Ancestors and survivors
Having calculated all primes up to a certain stage, we can trace back these primes for previous stages and see which of those have “survived” in the long run. All probable primes from stage 331 originated from one prime at stage 43. Recall that there were a total of 14 primes at stage 43, so the other 13 “petered out” along the way, including one branch that originated in stage 37 tenaciously keeping up until it succumbed at stage 95.
@@
Figure 3. Number of possible values for $y$ (for all we know). From $s=95$ onward, all primes $q$ emerge from one common ancestor at $s=43$, namely 1292942159746921794791923187781692727375711831825607985864285936838920812561. Naturally, the branches in the picture are not equally strong. For $s=331$ ($p=2221$), the following number of probable primes divide up on the nine given groups: 11477 + 9305 + 64706 + 123386 + 137475 + 101078 + 13648 + 77384 + 271928.
If one was to specify these values as $y_{0}$, $y_{1}$ and so on to attain fixed constants for the various possibilities, one might set $y_{0}$ as the minimum value $y_{min}$ as described above, and label each of the consecutive numbers $y_{1}$, $y_{2}$ … as the other smallest possible value(s) after a precalculated split, when it is “safe enough to assume” that those branches are stable.
@@
(from the first split at stage 44, $\lfloor y_{0}\cdot 191\#\rfloor\cdot 193+30$ and +88+88, $y_{1}\approx y_{0}+58/193\#$)
@@
The value $y_{6}$ is one that may be doubted more than $y_{0}$$y_{5}$ since it derives from the weakest branch of Figure 3 with only 9305 primes. But since 9305 >> 2221 (thus a stable branch), it can be expected to remain valid.
@@
The bifurcation tree as depicted in Figure 3 gets quite impressive when expanded. For example, tracing back the primes from stage 331 to stage 100 crystallizes out 47 primes (compared to originally 594) with an eventful history.
@@ 9. Broader utilization
It’s not without a little luck that the sequence as described exists. If we look for $y$’ such that $y’\cdot p\#$ rounded to the nearest integer is prime, we wouldn’t get far: at stage 23, one prime would be left: $\lfloor y’\cdot 83\#+\frac{1}{2}\rfloor=q’_{23,1}=206780313999369083332356327764879$. There’s no prime within $q’_{23,1}\cdot 89\pm 44$, so that sequence would be finite (or “mortal”) — even though the primes are smaller compared to $y=1.2541961$... ($q’_{s}=\{2,2,2,2,1,1,2,3,6,4,1,3,5,4,2,2,$ 2,4,5,5,2,2,1,0.}2,4,5,5,2,2,1,0.\} — and the ratio $\frac{y}{y’}=1.6198$... is close to the golden ratio:)
@@
$0<y<1$: this gives a sequence which is stronger than the original p#Y-sequence
($1<y<2$: original sequence)
$25<y<26$: stronger than the original sequence up to stage 166, weaker after stage 168
$1411<y<1412$: a weak sequence with only 883 primes at $s=168$, $p=997$
$3432<y<3433$: a very weak sequence, only 168 primes at $s=168$, $p=997$
$13948<y<13949$: another weak sequence
$201420<y<201421$: another one which has somewhat good chances to survive
$6007103<y<6007104$: rather weak
@@
$2<y<3$: extinct at stage 4
$3<y<4$: extinct at stage 7
$5<y<6$: extinct at stage 11
$16<y<17$: extinct at stage 14
$978<y<979$: extinct at stage 17
$6640<y<6641$: extinct at stage 23
$11456<y<11457$: extinct at stage 35
$160563<y<160564$: extinct at stage 38
$283257<y<283258$: extinct at stage 68
$1117230<y<1117231$: extinct at stage 78
$1594501<y<1594502$: extinct at stage 86
@@
Table 7. Data for $s\leq 330$, using BPSW-pseudoprimality for the primes $q$ for $s\geq 110$. nn* stands for the number of surviving branches, backtracked from the 810387 primes of stage 331. This number seems to be settled for $s\leq 130$ but will be subject to change for $s\geq 131$ by the time the calculation is taken further since many of them will get cancelled out, so this column only gives a momentary picture (as well as an upper bound). Nevertheless, the rate of decrease in nn* for larger $s$ may be of interest in itself.
$s$ $p$ $n$ $s$ $p$ $n$ nn* $s$ $p$ $n$ nn* $s$ $p$ $n$ nn*